OpenStudy (anonymous):

what two numbers equal 3.59

OpenStudy (anonymous):

1.45 and 2.14

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

1 + 159/200

OpenStudy (anonymous):

but the diffrence has to be 7 cent

OpenStudy (anonymous):

dude post the whole question

OpenStudy (anonymous):

two breakfest cearls, Kornies and Krispies togather cost 3.59 . one of the cost 7 cent more than the other. what is the difference in the price? which one cost more?

OpenStudy (anonymous):

x +x +7 = 3.59 solve for x

OpenStudy (anonymous):

then sub in x so one cereal costs x and the other cereal costs x +7

OpenStudy (anonymous):

how do i solve for x

OpenStudy (anonymous):

Use the formula x+x+0.07=3.59, where x represents the price of one cereal and x+1 represents the price of the other. Add up the x's to get 2x+0.07=3.59. Then isolate for the variable x, and you get the price of one cereal. Use the value you got for x, and substitute it into x+0.07 to get the price of the other.